# AI for Scientific Discovery is a Social Problem

### v3

@@ Abstract
Artificial intelligence promises to accelerate scientific discovery, yet its benefits remain unevenly distributed. While technical obstacles such as scarce data, fragmented standards, and unequal access to computation are significant, we argue that the primary barriers are social and institutional. Narratives that defer progress to speculative “AI scientists,” the undervaluing of data and infrastructure contributions, misaligned incentives, and gaps between domain experts and machine learning researchers all constrain impact. We highlight four interconnected challenges: community dysfunction, research priorities misaligned with upstream needs, data fragmentation, and infrastructure inequities. We argue that their roots lie in cultural and organizational practices. Addressing them requires not only technical innovation but also intentional community-building, cross-disciplinary education, shared benchmarks, and accessible infrastructure. We call for reframing AI for science as a collective social project, where sustainable collaboration and equitable participation are treated as prerequisites for technical progress.
@@ 1 Introduction
Artificial intelligence promises to accelerate scientific discovery, yet its benefits remain concentrated rather than democratized. While breakthrough applications like AlphaFold [49] and the Materials Project [44] demonstrate AI’s transformative potential, current development remains largely inaccessible to most researchers, requiring computational resources, specialized datasets, and interdisciplinary expertise that few institutions can match [72]. This concentration limits research diversity and undermines the collaborative foundations essential for robust scientific progress [26].
The challenge extends beyond resource scarcity to a fundamental mismatch between AI paradigms and scientific needs. Most AI development optimizes for predictive accuracy on large, homogeneous datasets, while science demands understanding from high-dimensional, low-sample-size data with complex relationships [10]. When AI tools prioritize automation over augmentation, they risk short-circuiting the human understanding essential for scientific progress [54], creating a production-progress paradox where increased output does not translate to deeper insight [50].
@@
This paper examines four critical barriers preventing AI democratization in science: community dysfunction that undermines collaboration, misaligned research priorities targeting narrow applications over upstream computational bottlenecks, data fragmentation due to incompatible standards [38], and infrastructure inequities concentrating power within privileged institutions [4]. While these manifest as technical challenges, their root causes are fundamentally social and institutional: including undervaluing data contributions, educational gaps preventing cross-disciplinary engagement, and absent community consensus on shared priorities. We then propose corresponding solutions: strengthening cross-disciplinary collaboration and education, structuring upstream challenges through shared benchmarks, standardizing scientific data practices, and building sustainable community-owned infrastructure. Our analysis demonstrates that democratizing AI for science requires treating it as a collective social project where equitable participation is a prerequisite for technical progress.
2 Barriers to Scientific Achievement
2.1 Barrier One: Community Dysfunction
The dysfunction within the AI for science community manifests through harmful narratives, misaligned incentives, and communication breakdowns that prevent effective collaboration. These reflect deeper cultural issues about how the community values different types of contributions, frames the purpose of AI in science, and structures interactions between disciplines.
@@ 2.1 Barrier One: Community Dysfunction
It has become popular to suggest that artificial general intelligence (AGI) is imminent and, that once it arrives, will solve the vast majority of our scientific problems [42]. Many efforts in the machine learning for science community have also focused on the development of so-called “AI scientists”, which have had results ranging from disappointing to actively harmful [27]. While these systems may eventually provide valuable support as co-pilot tools, particularly in retrieving accurate information and assisting with routine tasks, they are often framed as a future replacement for the human scientist.
This narrative is ultimately counterproductive to genuine scientific progress [15]. First, there is little evidence that AGI is imminent, let alone consensus on what the term means, and waiting for it risks delaying advances that could be achieved with current methods. Second, it devalues the expertise and contributions of human scientists, whose knowledge and creativity remain essential to discovery [45]. Third, it oversimplifies the inherent complexity of scientific practice, which depends not only on predictive accuracy but also on careful verification, contextualization, and theoretical integration [94]. Recent work by Luo et al.  [64] demonstrates that “AI scientists” are prone to failure modes, such as inappropriate benchmark selection, data leakage, metric misuse, and post-hoc selection bias, which are likely to be overlooked when evaluating only the final output. Finally, the vision of the fully autonomous “AI scientist” obscures the central purpose of science, which is not only to produce solutions but also to cultivate human understanding [24].
Rewarding Long-Lasting Impact
@@
Another counterproductive tendency within the current research ecosystem is the undervaluing of contributions to data and infrastructure in hiring, publicity, and tenure evaluations [34]. High-quality datasets, particularly those with well-curated metadata, often have far greater long-term impact than individual model contributions. Most models are rapidly superseded by marginally improved variants, and their influence diminishes within a short time frame. By contrast, datasets and infrastructure frequently underpin entire lines of research and remain useful for decades [73]. The effective half-life of data is therefore likely to be orders of magnitude longer than that of models. Recognizing and rewarding these contributions is essential for building a sustainable foundation for scientific progress with machine learning.
Technical Communication Gaps
Collaboration failures often stem from fundamental differences in how fields approach problem formulation and validation [55]. Domain scientists prioritize mechanistic understanding and experimental validation, while ML researchers focus on predictive performance and computational efficiency. Depending on the field, they may also rely on different tools and software practice, which further reinforces disciplinary boundaries. These different objectives lead to miscommunication about success criteria, acceptable trade-offs, and validation approaches.
@@
The terminology gap exacerbates these issues. Domain scientists may describe problems using field-specific concepts that don’t map directly to standard ML problem formulations. ML researchers may propose solutions that seem promising from a computational perspective but violate fundamental domain constraints or assumptions. For example, a machine learning model might achieve excellent accuracy in predicting rainfall patterns but do so by overfitting to seasonal cycles in the training data. Without incorporating physical constraints from climate dynamics, these predictions fail under shifting climate regimes and provide little scientific insight.
Insufficient Educational Materials
@@
A critical challenge lies in the lack of educational resources tailored for interdisciplinary research. Machine learning practitioners seeking to apply methods in biology, chemistry, and physics lack primers designed for technically mature but domain-naive audiences, while domain scientists frequently lack training in computational and ML methods. This bidirectional gap reflects the fact that most researchers trained in traditional sciences without computing requirements, often leading them to reinvent existing solutions [80]. Just as mathematics and statistics are fundamental to scientific training, machine learning should now be treated as essential for domain scientists. Without this foundation, collaborative efforts suffer as each group struggles to recognize what the other can contribute [13].
2.2 Barrier Two: Misaligned Research Priorities
@@
Several computational challenges represent similar opportunities for broad impact, for example:
  • Efficient PDE solvers: Generalizable solvers that work across different equations and report confidence when extrapolating beyond training data [86].
  • Multi-scale coupling: Linking atomistic simulations to continuum behavior for complex systems spanning multiple physical scales [25].
  • High-dimensional sampling: Underpinning uncertainty quantification and Bayesian inference across experimental sciences [14].
  • Quantum many-body methods: Tackling strongly correlated systems in condensed matter physics and quantum chemistry [79].
  • Fast materials property prediction: Enabling high-throughput screening for battery materials, catalysts, and semiconductors [90].
The challenge lies in identifying these upstream problems within specific research communities. Domain scientists often focus on their specialized applications without recognizing broader computational patterns. Machine learning researchers may lack the domain knowledge to identify which computational bottlenecks have the widest applicability.
@@
2.3 Barrier Three: Data Challenges
@@
Beyond architectural limitations, scientific data suffers from severe fragmentation that prevents collaborative model development. Research organizations typically manage over 100 distinct data sources, with 30% handling more than 1,000 sources [75]. For instance, a 2020 survey found that data scientists spend about 45% of their time on data preparation tasks, including loading and cleaning data [93]. Domain-specific formats, inconsistent metadata standards, and incompatible experimental protocols create artificial barriers that fragment knowledge and limit collaborative development.
@@
2.4 Barrier Four: Infrastructure Inequity
@@
3 Solutions and Implementation Strategies
If these barriers were addressed, the scientific community could achieve outcomes that remain out of reach today. Researchers across domains could build foundation models that integrate knowledge from biology, chemistry, physics and materials science rather than maintaining fragmented efforts. Experimental laboratories of all sizes could run rapid experiment–model loops that reduce iteration cycles from weeks to hours and accelerate discovery across disciplines. Scientists could rely on trustworthy AI co-pilots that highlight inconsistencies, suggest experiments and contextualize results. These systems would be grounded in robust data and reproducible infrastructure. Communities could rally around shared grand challenge benchmarks to drive progress on upstream computational problems. Most importantly, these advances would become accessible to a global community of researchers and would enable contributions from institutions and regions that are currently excluded by resource and infrastructure disparities.
3.1 Solution One: Strengthening Collaboration and Education Across Communities
@@
3.2 Solution Two: Structuring Upstream Challenges and Shared Benchmarks
@@
The research community needs systematic approaches to identify computational bottlenecks with broad applicability through dedicated conference sessions and cross-disciplinary workshops. Prize competitions demonstrate effective models for community-driven problem identification: the Vesuvius Challenge [2], DREAM Challenges [47] have engaged over 30,000 participants across sixty competitions producing 105+ publications, while DrivenData competitions [29] have provided \$4.8M+ in prizes across 245,000 submissions advancing climate, health, and education solutions. These competitions attract diverse expertise to problems individual research groups cannot tackle alone, using hybrid competition-collaboration models that prevent knowledge hoarding while maintaining competitive motivation.
@@
Benchmark datasets and evaluation metrics for upstream problems should focus on computational efficiency, accuracy across diverse problem instances, and uncertainty quantification rather than performance on narrow tasks [91]. For the scientific workforce to coalesce around such challenges, domain leaders must take responsibility for defining them and actively engaging their communities via easy to use open benchmark creation methods [81]. The most effective step individual researchers can take is to help articulate and participate in community benchmarks that frame upstream computational challenges as shared problems rather than siloed pursuits.
3.3 Solution Three: Standardizing and Curating Scientific Data for Broad Reuse
@@
Successful democratization of scientific data requires practical, widely adopted standards that balance usability with scientific integrity. Simple, universally compatible formats like CSV or Parquet often lower barriers to collaboration more effectively than complex, domain-specific schemas. Early bioinformatics efforts, such as MAGE-ML [83], illustrate the pitfalls of overly complex standards: despite technical flexibility, its complexity hindered interoperability, leading the community to adopt the simpler spreadsheet-based MAGE-TAB format [77, 76]. Similarly, attempts to harmonize nine Cancer Imaging Archive files required 329.5 hours of manual effort that could have been largely avoided with standardized schemas [20].
The Protein Data Bank (PDB) exemplifies large-scale, community-driven data democratization. Starting with seven structures in 1971, it now houses over 230,000 three-dimensional protein structures with 10 million daily downloads [8, 12]. Its success stems from standardized formats (PDB/mmCIF), expert curation, open access principles, and sustained funding. Similar approaches—minimal metadata standards, conversion tools for legacy formats, and validation pipelines—can enable other domains to achieve scalable reuse and collaborative development [73, 92, 39].
@@
The high-dimensional, low-sample-size problem requires architectural approaches that extend beyond current transformer models. Progress has been made in developing architectures tailored for scientific data. Graph neural networks show promise for molecular and materials applications by explicitly modeling spatial relationships, achieving 20% improvements in catalyst activity prediction over traditional descriptors [37, 36, 18]. Hierarchical attention mechanisms have been proposed to address the long-range interaction problem in biological systems, enabling analysis of gigapixel pathology images with spatial hierarchies spanning six orders of magnitude [16, 19, 41]. Physics-informed neural networks incorporate domain knowledge directly into model design, reducing simulation errors by 2-3 orders of magnitude in fluid dynamics applications while requiring 100x less training data [59, 89, 53]. State space models like Mamba [40] offer promising alternatives for handling long sequences common in scientific time series, achieving linear scaling with sequence length compared to quadratic complexity in transformers. Yet better inductive biases are still needed, and embedding them effectively into architectures remains non-trivial [85, 58]. Deep knowledge of both machine learning and the relevant domain is essential for success. Further, each of the research areas mentioned above lacks robust benchmarking and consequently has struggled to gain sustained traction [11].

### v4

@@ Abstract
Artificial intelligence (AI) is increasingly applied to scientific research, but its benefits remain unevenly distributed across communities and disciplines. While technical challenges such as limited data, fragmented standards, and unequal access to computational resources exist, social and institutional factors are often the primary constraints. Narratives emphasizing autonomous "AI scientists," under-recognition of data and infrastructure work, misaligned incentives, and gaps between domain experts and machine learning researchers all limit the impact of AI on scientific discovery. This paper highlights four interconnected challenges: community coordination, misalignment of research priorities with upstream needs, data fragmentation, and infrastructure inequities. We argue that addressing these challenges requires not only technical innovation but also intentional efforts in community-building, cross-disciplinary education, shared benchmarks, and accessible infrastructure. We call for reframing AI for science as a collective social project, where sustainable collaboration and equitable participation are treated as prerequisites for technical progress.
@@ 1 Introduction
In 2020, DeepMind’s AlphaFold solved a fifty-year grand challenge in biology: predicting protein structure from amino acid sequences with near-experimental accuracy [58]. The breakthrough earned a Nobel Prize and transformed structural biology [21]. Yet when researchers at under-resourced institutions attempted to apply similar methods to their own scientific problems, they encountered a different reality. They lacked the computational infrastructure to train comparable models, the curated datasets that made AlphaFold possible, and the interdisciplinary expertise to bridge machine learning and domain science. AlphaFold’s success demonstrated AI’s transformative potential for science. Its inaccessibility revealed a deeper challenge: AI tools may concentrate scientific capability rather than democratize it.
Recent breakthroughs illustrate AI’s expanding reach. AlphaFold3 now predicts interactions between proteins, DNA, RNA, and small molecules with at least 50% improvement over existing methods [1]. Google DeepMind’s GNoME discovered 2.2 million new crystal structures, including 52,000 novel lithium-ion conductors, with external researchers already synthesizing 736 predictions [79, 110]. Insilico Medicine’s ISM001-055 became the first fully AI-designed drug to reach Phase II trials with positive results, developed in under 18 months at approximately 10% of traditional costs [54]. BenevolentAI used AI to repurpose baricitinib for COVID-19, receiving FDA authorization within a year [109]. Climate modeling advanced with systems like NeuralGCM matching traditional models for 10-15 day forecasts with enormous computational savings [14, 69], NVIDIA’s StormCast super-resolving atmospheric data 1,000x faster using 3,000x less energy [87], and NASA-IBM’s Prithvi improving regional climate modeling resolution [82].
Almost all of these wonderful successes have a far less wonderful flip side. The benefits of these breakthroughs concentrate in domains with strong data infrastructure, well-coordinated research communities, and access to computational resources that most institutions cannot match [51]. The challenge extends beyond resource scarcity to fundamental questions about how scientific AI systems should be developed, shared, and governed. Without deliberate intervention, current development patterns risk deepening existing inequities in research capacity, limiting which questions get explored and who can contribute solutions.
What constitutes success in AI for science? We define scientific discovery as generating new mechanistic understanding about natural phenomena through systematic investigation, experimental validation, and integration with existing theoretical frameworks [31]. This contrasts with merely achieving predictive accuracy: a model that correctly forecasts protein binding affinity but provides no insight into which molecular interactions drive that binding represents a technical achievement but not necessarily scientific discovery. We define success along multiple dimensions: (1) predictive performance on well-defined tasks, (2) mechanistic insight revealing causal relationships and generalizable principles, (3) experimental validation confirming computational predictions in physical systems, (4) integration with domain knowledge advancing theoretical understanding, and (5) broad accessibility enabling researchers across institutions and geographies to build upon advances [64]. A truly successful AI application satisfies multiple criteria simultaneously, not just benchmark optimization.
A critical distinction exists between predictive performance and scientific understanding. As illustrated in Figure 1, a transformer-based model can achieve high accuracy predicting planetary trajectories without learning the underlying Newtonian mechanics [112]. General-purpose models optimized for prediction may not encode the mechanistic principles scientists seek to understand. Current commercial AI development disproportionately emphasizes literature synthesis, where large language models excel at processing existing text, over the experimental validation, iterative hypothesis testing, and mechanistic understanding that constitute scientific discovery’s core [59]. This risks conflating information retrieval with genuine discovery [64], potentially accelerating paper production while slowing actual scientific progress.
AI for science encompasses several distinct application categories, each with different technical requirements but facing common social and institutional barriers. Predictive modeling uses machine learning to forecast system behavior, such as protein structure prediction from amino acid sequences [58]. Simulation acceleration replaces expensive computational methods with faster surrogate models, enabling molecular dynamics and computational fluid dynamics at previously inaccessible scales [62, 39]. Design and optimization applies AI to iterative search in materials design and drug discovery [79, 54]. Literature synthesis and hypothesis generation leverages large language models to identify patterns across publications and suggest research directions [86, 96, 122]. Laboratory automation combines AI with robotics to accelerate experimental workflows [110]. Increasingly, researchers develop foundation models for science: large-scale pretrained models designed for scientific applications across chemistry, biology, and physics [84, 107]. These models aim to capture general scientific principles that can be fine-tuned for specific tasks with limited domain-specific data. While these applications have distinct technical characteristics, they share fundamental challenges rooted in how scientific communities organize, fund, and value different types of contributions.
@@
This paper examines four critical barriers preventing AI democratization in science: community dysfunction that undermines collaboration and perpetuates harmful narratives, research priorities targeting narrow applications over upstream computational bottlenecks, data fragmentation due to incompatible standards and social practices [44], and infrastructure inequities concentrating power within privileged institutions [3]. While these manifest as technical challenges, their root causes are fundamentally social and institutional: undervaluing data contributions, educational gaps preventing cross-disciplinary engagement, and absent community consensus on shared priorities. We propose corresponding solutions: strengthening cross-disciplinary collaboration and education, structuring upstream challenges through shared benchmarks, standardizing scientific data practices, and building sustainable community-owned infrastructure.
Although we distinguish social from technical barriers for clarity, the two are inseparable in practice. Every technical system embeds social assumptions about data collection, expertise, labor, incentives, and institutional constraints. The barriers we describe are interdependent rather than separate categories. We conclude that democratizing AI for science requires treating it as a collective social project where equitable participation is a prerequisite for technical progress.
Barriers to Scientific Achievement
Barrier One: Community Dysfunction
AI for science suffers from a cultural crisis. Harmful narratives about autonomous "AI scientists" coexist with systematic undervaluation of data work, while communication barriers between disciplines remain high. These aren’t merely coordination failures. They reflect deeper problems in how the community defines success, allocates recognition, and structures collaboration across fields.
@@
Recent commercial efforts by OpenAI, Anthropic, and Google DeepMind increasingly emphasize "AI scientists" capable of autonomous research [75, 6, 7]. OpenAI announced plans to develop an "intern-level research assistant" by September 2026 and a fully autonomous "legitimate AI researcher" by March 2028 [6], while Anthropic launched Claude for Life Sciences to "support the entire process, from early discovery through to translation and commercialization" [7]. These systems predominantly focus on literature search, summarization, and hypothesis generation from existing text (tasks where large language models demonstrate clear competence) while largely avoiding the experimental validation, iterative refinement, and mechanistic investigation that define scientific discovery.
This emphasis on literature synthesis creates several problems. First, it devalues experimental and computational research that generates new data rather than synthesizing existing publications. Second, it creates unrealistic expectations about AI’s current capabilities, leading to misallocation of resources toward tools that excel at information retrieval while neglecting those that could accelerate actual discovery [59]. Third, systems optimized for literature synthesis demonstrate serious failure modes when applied to research tasks, including inappropriate benchmark selection, data leakage, metric misuse, and post-hoc selection bias [76, 123].
The framing of autonomous "AI scientists" also perpetuates problematic attribution patterns. Historians and philosophers of science document how narratives of lone genius systematically erase the contributions of assistants, technical staff, and collaborative teams. These teams often include women and other marginalized groups whose labor proves essential but remains uncredited [41, 15]. When media headlines proclaim "AI scientist discovers X," they reproduce this problematic attribution by granting sole agency to the algorithm while obscuring the data curation, experimental validation, infrastructure development, and domain expertise that made the discovery possible. This erasure has material consequences: it devalues precisely the types of contributions (dataset construction, benchmark development, infrastructure maintenance) critical for democratizing AI in science.
Moreover, scientific creativity fundamentally depends on recognizing which problems are worth investigating. Callon describes this as the "struggles and negotiations" that define research agendas [19]. Current AI systems excel at optimization within well-defined problem spaces but lack capacity for problem formulation itself: identifying which questions matter, recognizing anomalous results that challenge existing frameworks, and judging when computational predictions warrant experimental follow-up. Framing AI as an autonomous scientist obscures this distinction between search execution and the creative, socially-embedded work of defining what to search for.
Conflicting Measures of Success and Impact
A fundamental tension exists between scientific discovery and AI optimization. Scientific success requires mechanistic insight: understanding why a prediction works, which variables matter, and how findings generalize beyond training conditions [31, 64]. AI success prioritizes predictive accuracy, optimization performance, and scaling behavior. A model that accurately predicts drug toxicity without identifying causal biochemical pathways may succeed as an AI system but fails as a scientific tool if it cannot guide rational design or explain unexpected failures. As illustrated in Figure 1, a transformer-based model can achieve high accuracy in predicting planetary trajectories without learning the underlying Newtonian mechanics [112]. This represents predictive performance without mechanistic understanding.
@@
This misalignment extends to how the research ecosystem rewards contributions. High-quality datasets, particularly those with well-curated metadata, often have far greater long-term impact than individual model contributions [38]. Most models are rapidly superseded by marginally improved variants and lose influence within months [25]. On the other hands, datasets and infrastructure frequently underpin entire lines of research and remain useful for decades [92]. Yet hiring, publicity, and tenure evaluations systematically undervalue data curation, infrastructure development, and benchmark creation in favor of novel algorithmic contributions. This creates perverse incentives: researchers pursue incremental model improvements that advance their careers while avoiding the foundational work that would enable broader progress.
Collaboration Gaps
Collaboration failures stem from fundamental differences in how fields approach problems [65]. Domain scientists prioritize mechanistic understanding and experimental validation. ML researchers focus on predictive performance and computational efficiency. Consider protein folding: a biologist asks which amino acid interactions drive stability, enabling rational mutation design. An ML researcher asks whether the model beats the benchmark, regardless of interpretability. In climate science, domain experts need models that respect physical constraints like conservation of mass and energy, whereas ML practitioners might propose purely data-driven approaches that achieve better short-term predictions but fail catastrophically under novel conditions.
@@
The terminology gap exacerbates these issues. Domain scientists describe problems using field-specific concepts that don’t map to standard ML formulations. ML researchers propose solutions that seem computationally promising but violate fundamental domain constraints. A machine learning model might achieve excellent accuracy predicting rainfall patterns by overfitting to seasonal cycles in training data. Without physical constraints from climate dynamics, these predictions fail under shifting climate regimes and provide little scientific insight.
@@
Educational gaps compound collaboration challenges. Machine learning practitioners seeking to apply methods in biology, chemistry, and physics lack primers designed for technically mature but domain-naive audiences. Domain scientists frequently lack training in computational and ML methods, reflecting the fact that most trained in traditional sciences without computing requirements. This often leads them to reinvent existing solutions [103]. Just as mathematics and statistics are fundamental to scientific training, machine learning should now be treated as essential for domain scientists. Without this foundation, collaborative efforts suffer as each group struggles to recognize what the other can contribute [16].
Barrier Two: Misaligned Research Priorities
@@
To move beyond narrow applications, the community must target upstream computational bottlenecks. Initiatives such as the Fundamental Development Gap Map [28] have begun cataloging these hurdles: specific mathematical or algorithmic barriers that, if cleared, would unlock progress across dozens of downstream disciplines. Some example priority challenges are:
  • Generalizable PDE Neural Operators. Current simulations for weather, plasma fusion, and aerodynamics depend on expensive numerical PDE solvers. The upstream challenge is to develop physics foundation models that learn operator mappings rather than individual solutions. The main unsolved hurdle is robust generalization across regimes, such as moving from laminar to turbulent flow or from present to future climate states. Progress here would transform our ability to model complex systems, although true regime extrapolation remains an open problem [62, 14, 107].
  • The Inverse Problem in High Dimensions. Forward modeling that predicts properties from structure is now common, but the inverse task of recovering structure from desired properties is still the primary barrier in materials science, drug discovery, and several areas of astrophysics. A broadly applicable probabilistic framework that can invert complex, nonlinear physical models would shift many disciplines from empirical screening to rational design. Existing approaches solve special cases but do not yet provide a general solution [117, 79].
  • The Sampling Problem in Rare Events. Many important scientific phenomena, such as protein folding transitions, extreme weather, and chemical reaction barriers, occur in the low-probability tails of high-dimensional distributions. Standard AI training procedures focus on average performance and rarely capture these events. Methods for efficient importance sampling in high dimensions are still immature. Improving rare-event sampling would accelerate discovery in stochastic systems and strengthen risk analysis for safety-critical applications [17, 85].
  • Multi-Scale Biological Coupling. Biological systems span a hierarchy of spatial and temporal scales, from angstrom-level molecular dynamics to millimeter-scale tissue mechanics. Integrating these regimes in a unified model remains unsolved. Current virtual-cell initiatives often rely on single-cell transcriptomic models that do not incorporate mechanical or spatial constraints. Solving the multi-scale coupling problem is a prerequisite for realistic biological digital twins, although partial frameworks exist at individual scales [61, 18].
  • Computing Strongly Correlated Quantum Systems. Accurately modeling strongly correlated electrons remains one of the central computational barriers in quantum chemistry and condensed matter physics. Challenges such as the fermion sign problem limit the scalability of Monte Carlo and tensor-network methods, and no general-purpose approach avoids exponential growth in complexity. AI-driven variational models offer promising approximations for specific systems, but they do not resolve the broader difficulty of representing highly entangled many-body states in large geometries. Progress here would enable predictive design of quantum materials, including unconventional superconductors and advanced catalytic systems [101].
@@
Barrier Three: Data Challenges
@@
Beyond architectural limitations, scientific data suffers from severe fragmentation that prevents collaborative model development. Research organizations typically manage over 100 distinct data sources, with 30% handling more than 1,000 sources [95]. For instance, a 2020 survey found that data scientists spend about 45% of their time on data preparation tasks, including loading and cleaning data [121]. Domain-specific formats, inconsistent metadata standards, and incompatible experimental protocols create artificial barriers that fragment knowledge and limit collaborative development. Recent research on machine learning data practices reveals fundamental challenges: ML researchers struggle to apply standard data curation principles due to shared terms with non-shared meanings between fields, high interpretative flexibility in adapting concepts, the depth of data curation expertise required, and unclear scoping of documentation responsibilities [50].
@@
Barrier Four: Infrastructure Inequity
@@
Solutions and Implementation Strategies
If these social and technical obstacles were meaningfully addressed, the scientific community could begin to unlock more of AI’s long-term promise, though not without continued risk, investment, and coordination. For example, researchers from biology, chemistry, physics, and materials science might more systematically pool data and methods to build shared, cross-domain models, rather than working in isolated silos [84]. Laboratories of varying sizes could deploy tighter experiment-model feedback loops that compress cycles from weeks to days, as demonstrated by autonomous laboratories that have successfully synthesized dozens of new materials in continuous operation [110]. With better standards for provenance, metadata, and reproducibility, trusted AI assistants could help scientists spot anomalies, suggest hypotheses, and contextualize findings, while humans remain central to interpretation and validation. At the same time, communities could rally around common benchmarks and shared “grand challenge” problems to coordinate collective progress across disciplines. Importantly, more equitable infrastructure, whether open datasets, low-cost compute platforms, or regional capacity-building, could broaden participation, making these capabilities accessible to researchers in under-resourced institutions or geographies.
Solution One: Strengthening Collaboration and Education Across Communities
@@
Case Study: Interdisciplinary Training Models
The Eric and Wendy Schmidt AI in Science Postdoctoral Fellowship, a program of Schmidt Sciences, demonstrates effective approaches to building interdisciplinary capacity. The program annually supports approximately 160 postdoctoral fellows across nine universities globally [102]. Fellows receive dual mentorship combining domain science expertise with AI technical skills, recognizing that effective AI for science requires specialized training rather than superficial collaboration between separate teams. The program structures cohort-based learning where fellows develop shared vocabulary across disciplines, provides access to computational resources, and creates explicit career pathways that reward bridge-building contributions. Early outcomes suggest that dedicated interdisciplinary roles, when properly supported institutionally, can overcome collaboration barriers more effectively than ad hoc partnerships between isolated research groups.
Solution Two: Structuring Upstream Challenges and Shared Benchmarks
@@
The research community needs systematic approaches to identify computational bottlenecks with broad applicability through dedicated conference sessions and cross-disciplinary workshops. Several machine learning communities have already begun addressing data and benchmark gaps. For example, the NeurIPS Datasets and Benchmarks Track has established expectations around documentation, evaluation protocols, and dataset governance. However, these efforts remain largely internal to the ML research ecosystem and are not yet integrated with the domain-specific validation pipelines, experimental constraints, and long-term data stewardship required for scientific discovery.
Meaningful AI-for-science benchmarks require sustained coordination between domain scientists, experimental facilities, and ML researchers, not only standardized ML evaluation formats. To that end, prize competitions demonstrate effective models for community-driven problem identification: the Vesuvius Challenge [114], DREAM Challenges [55] have engaged over 30,000 participants across sixty competitions producing 105+ publications, while DrivenData competitions [33] have provided \$4.8M+ in prizes across 245,000 submissions advancing climate, health, and education solutions. These competitions attract diverse expertise to problems individual research groups cannot tackle alone, using hybrid competition-collaboration models that prevent knowledge hoarding while maintaining competitive motivation.
@@
Benchmark datasets and evaluation metrics for upstream problems should focus on computational efficiency, accuracy across diverse problem instances, and uncertainty quantification rather than performance on narrow tasks [118]. Recent work advocates for comprehensive benchmark suites rather than single leaderboard scores, allowing better understanding of trade-offs between AI models without hiding potential harms within aggregate metrics [4]. For the scientific workforce to coalesce around such challenges, domain leaders must take responsibility for defining them and actively engaging their communities via easy to use open benchmark creation methods [105]. The most effective step individual researchers can take is to help articulate and participate in community benchmarks that frame upstream computational challenges as shared problems rather than siloed pursuits.
Case Study: CASP and Upstream Problem Structuring
The Critical Assessment of protein Structure Prediction (CASP) exemplifies how community coordination around upstream computational challenges enables transformative breakthroughs. Launched in 1994 by John Moult and Krzysztof Fidelis, CASP established biennial blind prediction competitions with standardized evaluation protocols [81]. Participating groups submit structure predictions before experimental structures are released, ensuring objective assessment free from confirmation bias.
CASP’s success stems from social infrastructure as much as technical design. The organizers built institutional norms around open participation, rigorous evaluation, and shared problem definition that persisted across decades and leadership transitions. Target selection protocols ensured predictions remained truly blind, while multiple assessment methods provided consensus validation. The competition format created both collaborative knowledge-sharing (through post-assessment workshops and publications) and competitive motivation to advance methods. Sustained NIH funding from 1994 to 2025 enabled continuity, though recent funding challenges underscore infrastructure fragility when government support wavers.
This groundwork proved critical for AlphaFold’s breakthrough at CASP14 in 2020 [58]. DeepMind’s technical innovation succeeded precisely because CASP had established clear success metrics, assembled challenging test cases, and created a community ready to validate and build upon advances. Without CASP’s two decades of problem structuring and community coordination, even sophisticated neural architectures would have struggled to demonstrate meaningful progress on protein structure prediction. The case illustrates how upstream problem structuring through sustained community investment can multiply downstream impact compared to fragmented efforts on narrow applications.
Solution Three: Standardizing and Curating Scientific Data for Broad Reuse
@@
Successful democratization of scientific data requires practical, widely adopted standards that balance usability with scientific integrity. Simple, universally compatible formats like CSV or Parquet often lower barriers to collaboration more effectively than complex, domain-specific schemas. Early bioinformatics efforts, such as MAGE-ML [108], illustrate the pitfalls of overly complex standards: despite technical flexibility, its complexity hindered interoperability, leading the community to adopt the simpler spreadsheet-based MAGE-TAB format [99, 98]. As noted earlier regarding the Cancer Imaging Archive [26], the lack of standardized schemas necessitates hundreds of hours of manual labor—effort that could be obviated by adopting simple, consistent formats.
The Protein Data Bank (PDB) exemplifies large-scale, community-driven data democratization. Starting with seven structures in 1971, it now houses over 230,000 three-dimensional protein structures with 10 million daily downloads [10, 13]. Its success stems from standardized formats (PDB/mmCIF), expert curation, open access principles, and sustained funding. Similar approaches—minimal metadata standards, conversion tools for legacy formats, and validation pipelines—can enable other domains to achieve scalable reuse and collaborative development [92, 119, 45]. The FAIR principles (Findable, Accessible, Interoperable, and Reusable) are increasingly being applied not just to data but also to AI models and software, though implementation remains challenging and requires community consensus [52]. Tools like AIDRIN (AI Data Readiness Inspector) now provide comprehensive frameworks for assessing data readiness including FAIR compliance, quality, understandability, value, fairness, and bias metrics [47]. The 2021-2024 FAIR Data Spaces project in Germany laid foundations for common data spaces from organizational, legal, and technical perspectives [40].
@@
The high-dimensional, low-sample-size problem requires architectural approaches that extend beyond current transformer models. Progress has been made in developing architectures tailored for scientific data. Graph neural networks show promise for molecular and materials applications by explicitly modeling spatial relationships, achieving 20% improvements in catalyst activity prediction over traditional descriptors [43, 42, 23]. Hierarchical attention mechanisms have been proposed to address the long-range interaction problem in biological systems, enabling analysis of gigapixel pathology images with spatial hierarchies spanning six orders of magnitude [20, 24, 48]. Physics-informed neural networks incorporate domain knowledge directly into model design, reducing simulation errors by 2-3 orders of magnitude in fluid dynamics applications while requiring 100x less training data [70, 116, 63]. State space models like Mamba [46] offer promising alternatives for handling long sequences common in scientific time series, achieving linear scaling with sequence length compared to quadratic complexity in transformers. Yet architectures still need better inductive biases about data structure and relationships, and embedding such assumptions effectively remains non-trivial [112, 68]. Deep knowledge of both machine learning and the relevant domain is essential for success. Further, each of the research areas mentioned above lacks robust benchmarking and consequently has struggled to gain sustained traction [12].